OpenStudy (starryknight):

please help..... explain in steps Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 5x2 + 20x - 120 = 0 answer: (x + 2)2 = 28 (x + 5)2 = -21 (x + 2)2 = -28 (x + 4)2 = 21

OpenStudy (kirbykirby):

If your original quadratic is \[ ax^2+bx+c=0\], then you first divide through by \(a\) to get rid of the leading coefficient. Then you want to add and subtract the term: \(\large \left(\frac{b}{2} \right)^2\) \(5x^2+20x-120 =0 \), divide by 5 \(x^2+4x-24 = 0\) Now add and subtract \(\large \left( \frac{4}{2}\right)^2=4\) \[ \underbrace{x^2+4x+ 4}-4-24=0\\ \underbrace{(x+2)^2}-28=0\\ (x+2)^2=28\]

OpenStudy (starryknight):

thanx :)

OpenStudy (kirbykirby):

=]